Lana is the oldest of four sisters. Her youngest sister is half her age. The other two sisters are twins 2 years younger than Lana. When the four sisters' ages are added together, the result is 45. If the youngest sister is n years old, what is n?
Found 2 solutions by greenestamps, mananth:
Answer by greenestamps(13200)



Let n be the age of the youngest.

Then Lana's age is 2n.

Then the age of each of the twins is 2n-2.

The sum of all four ages is 45:

(n)+(2n)+(2n-2)+(2n-2) = 45
7n-4 = 45
7n = 49
n = 7

ANSWER: The age of the youngest is 7

CHECK:
youngest: n = 7
Lana: 2n = 14
1st twin: 2n-2 = 12
2nd twin: 2n-2 = 12
7+14+12+12=45


Answer by mananth(16946)

Lana is the oldest of four sisters. Her youngest sister is half her age. The other two sisters are twins 2 years younger than Lana. When the four sisters' ages are added together, the result is 45. If the youngest sister is n years old, what is n?


Let youngest sister be n years old
Lana is twice the youngest sister's age. =2n
The other two sisters are twins 2 years younger than Lana.
The age of the twins is (2n-2) years
When the four sisters' ages are added together, the result is 45.
n+2n +2*(2n-2)=45
n+2n +4n-4=45

7n -4=45
7n = 49
n =49/7 = 7



14+12+12+7 =45